OpenStudy (anonymous):

Find an angle which is 1/2 its complement

OpenStudy (anonymous):

faye!!! x+.5x=90

OpenStudy (anonymous):

60 ung ans

OpenStudy (anonymous):

panong 60?

OpenStudy (anonymous):

edi magiging 6x=90 x=15?

OpenStudy (anonymous):

90/1.5x

OpenStudy (anonymous):

The sum of an angle and its complement is 90degrees\[\theta + \frac{\theta}{2} = 90\]Solve for theta.

OpenStudy (anonymous):

kasi nag x+.5x

OpenStudy (anonymous):

*nga ndi nag

OpenStudy (anonymous):

ah cgecge tnx